OpenStudy (anonymous):

A chairman is biased so that he selects his relatives for a job 3 times as likely as others. If there are 3 posts for a job. Find the probability distribution for selection of persons other than their relatives. If the chairman is biased than which value of life will be demolished?

OpenStudy (anonymous):

challenging question

OpenStudy (anonymous):

.25 i guess

OpenStudy (anonymous):

do we need a relative quantification of the number of relatives in the probability pool?

OpenStudy (anonymous):

For 0 relatives, it is 100% For 3 relatives, it is 0%

OpenStudy (kropot72):

Let the probability of selecting an other person = x Then probability of selecting a relative = 3x \[P(0\ others)=\left(\begin{matrix}3 \\ 0\end{matrix}\right)x ^{0}(1-x)^{3}=(1-x)^{3}\] \[P(3\ relatives)=\left(\begin{matrix}3 \\ 3\end{matrix}\right)(3x)^{3}(1-3x)^{0}=(3x)^{3}\] \[P(0\ others)+P(3\ relatives)=1\] \[(1-x)^{3}+(3x)^{3}=1\ .......................(1)\] The non-zero, positive solution of equation (1) is x=0.2868 Using the value of 0.2868 for the probability of selecting a person other than a relative the binomial distribution can be used to find P(0 others), P(1 other), P(2 others) and P(3 others).

OpenStudy (anonymous):

also how ur 0.2868 comes, plse explain

OpenStudy (anonymous):

If you are doing Bernoulli trial, I would suggest 1:3 converted to \(p=1/4,q=3/4\)

OpenStudy (anonymous):

\[ \Pr(X=0) = \binom{3}{0}\left(\frac{1}{4}\right)^0\left(\frac{3}{4}\right)^3 = \frac{27}{64}\\ \Pr(X=1) = \binom{3}{1}\left(\frac{1}{4}\right)^1\left(\frac{3}{4}\right)^2 = \frac{27}{64}\\ \Pr(X=2) = \binom{3}{2}\left(\frac{1}{4}\right)^2\left(\frac{3}{4}\right)^1 = \frac{9}{64}\\ \Pr(X=3) = \binom{3}{3}\left(\frac{1}{4}\right)^3\left(\frac{3}{4}\right)^0 = \frac{1}{64} \]

OpenStudy (kropot72):

But if you use 0.25 for the probability of choosing a relative and 0.75 for the probability of choosing other than a relative the sum of Pr(X = 0) + Po(X = 3) does not equal 1. Also how did you find the value p = 1/4 ?

OpenStudy (anonymous):

To find \(p\) I used a system of equations. \[ p = 3q \\ p = 1-q \]

OpenStudy (anonymous):

The sum \(\Pr(X=0) + \Pr(X=3)\) should not equal \(1\), because there is nonzero probability that \(X=1,X=2\)

OpenStudy (kropot72):

My bad. The sum is Po(X = 0) + Pr(X = 3) where Po is the probability of other than a relative being selected and Pr is the probability of a relative being selected. If Po = 1/4 and Pr = 3/4 this sum does not equal 1. Therefore your values are not a valid solution.

OpenStudy (anonymous):

When you say 'other than a relative being select' and 'a relative being selected'... You're not talking about mutually exclusive events, so due to double counting, they should be greater than 1 when added.

OpenStudy (anonymous):

In my model, 'other than a relative being select' and 'a relative being selected' would be represented by \(\Pr(X\gt 0)\) and \(\Pr(X\lt 3)\) respectively.

OpenStudy (kropot72):

I agree that my reasoning is flawed :(

OpenStudy (anonymous):

@wio the concept of 4 is not understood to me plse explain

OpenStudy (anonymous):

Is this advanced probability? College level?

OpenStudy (kropot72):

Let the probability that a person other than a relative is selected = x Then the probability that a relative is selected = 3x The sum of the above probabilities must equal one. Therefore x + 3x = 1 4x = 1 x = 1/4 = 0.25 Then use the formula for the binomial distribution with n = 3 and p = 0.25 to find P(X = 0), P(X = 1), P(X = 2), P(X = 3).

OpenStudy (anonymous):

@msingh Can you be more specific about what you don't understand? Do you not understand the Binomial Distribution? Do you not understand how I found \(p\) and \(q\)?

OpenStudy (anonymous):

@wio i know binomial theorem and also understood the concept of 4 thank u once again